OpenStudy (anonymous):

Anthony wants to make 6 gallons of a 34% alcohol solution by mixing together 24% alcohol solution and a 64% alcohol solution. How many of each solution must he use.

OpenStudy (kropot72):

Let a be the required amount of 24% alcohol solution. Let b be the required amount of 64% alcohol solution. Now we can write: a + b = 6 .......................(1) The amount of alcohol in 6 gallons of 34% alcohol solution is given by: 6 * 0.34 = 2.04 ..............(2) Therefore we can write another equation based on the alcohol content in the three different alcohol solutions: 0.24a + 0.64b = 2.04 ......(3) Equations (1) and (3) are a pair of simultaneous equations which can be solved to find the values of a and b.

OpenStudy (anonymous):

Thank you for the help

OpenStudy (anonymous):

Haha satellite73

OpenStudy (kropot72):

You're welcome :)